OpenStudy (anonymous):

please help!! I have absolutely no idea how to do this! Write the expression as the sine, cosine, or tangent of an angle. sin 9x cos x - cos 9x sin x

OpenStudy (anonymous):

the options are -sin 10x -cos 8x -sin 8x -cos 10x

OpenStudy (anonymous):

look up sum formulas for sin and cos

OpenStudy (anonymous):

like the identities?

OpenStudy (anonymous):

oh and i just noticed that i made the options look like negatives but none of them are anyways sin x=1/csc x cos x=1/sec x ^^ that?

OpenStudy (anonymous):

no... have a look at the link i sent

OpenStudy (anonymous):

ohh just saw the llink my bad.. ok so where do i go from there?

OpenStudy (anonymous):

have a look and see if any match what you have going on and go from there

OpenStudy (anonymous):

sin (a-b)=sin a cos b-cos a sin b?

OpenStudy (anonymous):

that's what it looks like to me

OpenStudy (anonymous):

so it would be sin(9x-x) so sin 8x right?

OpenStudy (anonymous):

that's what it looks like to me

OpenStudy (anonymous):

yayy thank you soo much!!

OpenStudy (anonymous):

you're welcome!!!

OpenStudy (anonymous):

how can I give you a medal? I am new to this site..

OpenStudy (anonymous):

just click on best response, but it's not necessary. just glad I could help

OpenStudy (anonymous):

haha did it anyways thank you again.. Im finally starting to get it a bit more

OpenStudy (anonymous):

awesome! keep at it!